OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2).

OpenStudy (anonymous):

The equation form is y=ax+b.

OpenStudy (lgbasallote):

use the formula \[y - y_1 = m(x-x_1)\] y1 is the y coordinate x1 is the x coordinate and m is the slope just substitute into the formula. it's very straightforward

OpenStudy (anonymous):

y-(-2)=3(x-1)

OpenStudy (anonymous):

y+2=3x-3

OpenStudy (anonymous):

y=3x-3-2

OpenStudy (anonymous):

y=3x-5

OpenStudy (anonymous):

oh okay, I understand. so what I had to do was plug the numbers in.

OpenStudy (anonymous):

Yes, just plug the numbers.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You are welcome